Capt. Hooke is on a treasure hunt. His GPS shows him that he is 25 m away from what he is looking for. He walks 25 m towards west. The GPS compass now tells him that the treasure is due south from where he is standing. How far south does he need to go to discover it?


1
Expert's answer
2021-05-24T15:19:20-0400

Capt. Hooke goes 25m south

let's say he needed to go x m south

Hence total distance from start


"=\sqrt{25^2+X^2}"


25m away it was from starting hence



"\sqrt{25^2+X^2}=25\\\implies X=0"

Therefore, Capt. Hooke does not need to walk any more distance. He just needs to rotate 90 degrees on the left.